The formula for an anti-itch lotion includes 15 ml of calamine in a total of 60 ml of solution. How much calamine is needed to make 90 ml of lotion?
Answer by star68(2)

15ml of calamine to 60ml of solution
60/6= 10ml
15/6=2.5ml
This means that there are 2.5ml of calamine per 10ml of anti-itch lotion, thus for a 90ml of lotion you multiply 9 to both the solution and calamine=
10*9=90ml
2.5*9=22.5ml
Therefore, there are 22.5ml of calamine in a 90ml anti-itch lotion